OpenStudy (fools101):

Exponential Equations... The population of a small coastal resort town, currently 2500, grows at a rate of 1.5% per year. a) Write an exponential equation to model this information. b) Find the number of years for the population to reach 5,000. Exponential Equations... The population of a small coastal resort town, currently 2500, grows at a rate of 1.5% per year. a) Write an exponential equation to model this information. b) Find the number of years for the population to reach 5,000. @Mathematics

OpenStudy (fools101):

Please Help!!!

OpenStudy (wasiqss):

w8 solvin

OpenStudy (fools101):

I DON"T KNOW THAT ALL IT SHOWS! Exponential Functions The population of a small coastal resort town, currently 2500, grows at a rate of 1.5% per year. a) Write an exponential equation to model this information. b) Find the number of years for the population to reach 5,000.

OpenStudy (wasiqss):

gal ur question r so confusing

OpenStudy (fools101):

I know -_- but that what it said! *Sigh**

OpenStudy (wasiqss):

(1+.015)^ number of years then multiply by 2500

OpenStudy (wasiqss):

the answer to b part is 46.55 or 47 to nearest whole num

OpenStudy (wasiqss):

because 5000=2500(1.015)^n then 2=(.1.015)^n take ln both sides ln2/ln1.015= n answer will be yieled

OpenStudy (wasiqss):

dum dum :P

OpenStudy (fools101):

Thankssss -_-!